OpenStudy (babyslapmafro):

Find the equation of the line tangent to the graph of tanx at x=0.

OpenStudy (babyslapmafro):

f'(x)=sec^2(x)

OpenStudy (babyslapmafro):

y-tanx=sec^2(x)(x-0)

OpenStudy (babyslapmafro):

y=xsec^2(x)+tanx

OpenStudy (babyslapmafro):

Where am I going wrong?

OpenStudy (anonymous):

y=xsec^2(x)+tanx needs more pie